# Supersymmetric scale-separated AdS3 vacua of type IIB

### v1

@@ Abstract
I construct supersymmetric, parametrically scale-separated AdS3 vacua of type IIB string theory. These arise as compactifications with orientifold planes on specific seven-dimensional solvmanifolds admitting co-closed $G_{2}$-structures, preserving minimal supersymmetry. There are solutions that include either one set or four sets of intersecting O5-planes in the smeared approximation, and parametric scale separation can be achieved by tuning unbounded fluxes to infinity. Additionally, the putative holographic field theory operators that are dual to the lightest scalars in the gravitational theory have integer conformal dimensions at tree level, aligning with other scale-separated models of type II string theory.
@@ 1 Introduction
where $L_{\mathrm{KK}}$ and $L_{\mathrm{AdS}}$ are the KK and AdS radius respectively. Some recent progress has provided strong arguments against scale separation in specific settings. For instance, refs. Cribiori:2022trc ; Bobev:2023dwx ; Perlmutter:2024noo indicate that scale separation for compactifications of several theories with extended supersymmetry is unattainable. Moreover, it has been argued in ref. Gautason:2015tig that for compactifications without sources, scale separation cannot occur for isotropic manifolds of positive scalar curvature, more precisely for when the KK scale decouples from the scalar curvature, and due to the lack of concrete examples it was conjectured not to happen in ref. Collins:2022nux . All these findings support the strong version of the Anti-de Sitter distance conjecture Lust:2019zwm , implying that no supersymmetric scale-separated vacua exist. However, the minimally supersymmetric theories with sources are not constrained yet by the arguments of refs. Cribiori:2022trc ; Bobev:2023dwx ; Perlmutter:2024noo ; Gautason:2015tig ; Collins:2022nux , and indeed, there are a few scale-separated models in string theory, therefore violating the Anti-de Sitter distance conjecture Lust:2019zwm . Nevertheless, their validity is questioned and is an active research topic in the context of the swampland program. For a recent review on this issue, see ref. Coudarchet:2023mfs .
@@
The current state of the art of supersymmetric scale-separated vacua can be summarised as a type IIA vs. IIB dichotomy: in type IIA, there are scale-separated vacua where all moduli are stabilised using fluxes and orientifold planes, hence they are referred to as classical vacua. The separation of scales is achieved by tuning unbounded, integer-quantised fluxes to infinity. Examples in four dimensions are compactifications on manifolds with SU(3)-structure, usually toroidal orbifolds, e.g. the DGKT solutions in massive IIA DeWolfe:2005uu ; Camara:2005dc and analogue solutions in massless IIA Cribiori:2021djm ; Carrasco:2023hta , of which some can be lifted to solutions of M-theory which appear to have weak $G_{2}$-structure VanHemelryck:2024bas . Examples in three dimensions are compactifications on manifolds with $G_{2}$-holonomy, as considered in Farakos:2020phe ; VanHemelryck:2022ynr . [^1] See also refs. Farakos:2023nms ; Farakos:2023wps ; Tringas:2023vzn for variations on these setups. Although the holographic dual field theories to these models are not known, refs. Conlon:2021cjk ; Apers:2022tfm ; Quirant:2022fpn found, quite surprisingly, that the mass spectrum of the scalars in the four-dimensional scale-separated vacua is such that the putative holographic dual operators in the field theory have integer conformal dimensions. [^2] This feature is expected to break down by small one-loop corrections, see ref. Plauschinn:2022ztd . This feature does not appear in the three-dimensional vacua of ref. Farakos:2020phe , as shown in ref. Apers:2022zjx .
@@
The situation in type IIB is quite different. Non-perturbative effects are usually necessary for full moduli stabilisation and obtaining scale separation. Examples are four-dimensional Calabi-Yau compactifications with three-form fluxes and a gaugino condensate or Euclidean D3-instantons, such as in the KKLT and LVS scenarios Kachru:2003aw ; Conlon:2005ki ; Balasubramanian:2005zx . Recent progress has been made in finding solutions with a large hierarchy of scales (i.e. finding a small flux superpotential), see for example refs. Demirtas:2019sip ; Demirtas:2020ffz ; Demirtas:2021nlu ; Demirtas:2021ote ; McAllister:2024lnt , however, these scenarios have been criticised on different fronts as well, see e.g. refs. Bena:2020xrh ; Bena:2018fqc ; Gao:2020xqh ; Gao:2022fdi ; Gao:2022uop ; Lust:2022lfc for the most recent criticism. So far, no supersymmetric and scale-separated vacua of type IIB have been found where only classical ingredients stabilise all moduli, while maintaining weak coupling and large volume. [^3] The 4d vacua of refs. Petrini:2013ika ; Caviezel:2009tu are not well-controlled, as some one-cycles are becoming small in string units, see ref. Cribiori:2021djm .
@@
However, this paper presents the first such examples. These are realised by compactifying type IIB string theory on seven-dimensional manifolds with a co-closed $G_{2}$-structure. Compactifications on such backgrounds have been considered before in refs. Emelin:2021gzx ; Passias:2020ubv ; VanHemelryck:2022ynr . Here, two specific orbifolds of a particular seven-dimensional solvmanifold are used. This solvmanifold has been described in the literature before, see e.g. refs. Andriot:2015sia ; Manero2020315 . After orientifolding these specific spaces with an O5-involution, one obtains one class of manifolds with four (intersecting) sets of (smeared) O5-planes and another class with only one set of O5-planes. The former is reminiscent of the type IIA vacua with 4 sets of (intersecting) O6-planes, whereas the latter case can be regarded as the supersymmetric cousin of a scale-separated, but non-supersymmetric type IIB vacuum found in ref. Arboleya:2024vnp , which are related through skew-whiffing. The analysis here uses the tools of refs. Emelin:2021gzx ; Passias:2020ubv ; VanHemelryck:2022ynr , but is different in the sense that in those references, models with seven (intersecting) sets of O5-planes were considered instead of fewer. Particularly in ref. Emelin:2021gzx , it was argued that scale separation can be achieved in their setup if one includes additional D5-branes to cancel some of the tadpoles, and if one can make the curvature small compared to the compactification radius. However, the latter clashes with geometric flux quantisation when the manifold is isotropic, and no further explicit models were considered either. The setups of this paper circumvent these issues and do not need the inclusion of D5-branes. Additionally, when computing the masses of the scalars in the untwisted sector, one obtains that the conformal dimensions of the corresponding operators in the putative holographic dual theory, are all integer, sharing this feature with their four-dimensional counterparts.
This paper is organised as follows: section 2 discusses the seven-dimensional solvmanifold and the different orbifold groups that it is quotiented by, such that an orientifold involution produces the right O5-plane content for the type IIB solutions of interest. Section 3 contains a brief review of type IIB compactifications on manifolds with $G_{2}$-structure, and discusses the specific compactifications on the manifolds introduced in the previous section and their scale separation feature. In section 4, the mass spectrum and conformal dimensions are discussed. Finally, the results are summarised in section 5.
2 A seven-dimensional solvmanifold, orbifolded and orientifolded
@@
Before considering the details of three-dimensional flux compactifications of type IIB string theory, it is useful to focus on the specific compact space that is used later in section 3. In particular, solvmanifolds are considered throughout this paper, which are group manifolds arising from a solvable group. There exists a variety of such solvmanifolds across different dimensions, amongst which many have been used for compactifications before, for instance in refs. Grana:2006kf ; Andriot:2010ju ; Danielsson:2011au ; Grana:2013ila ; Andriot:2015sia ; Andriot:2022way ; Andriot:2022yyj . Here, a very specific seven-dimensional solvmanifold is considered. It arises as a twisted torus, denoted by $\tilde{\mathbb{T}}^{7}$, for which the metric in the vielbein parametrisation is
@@
where the $e^{i}=e^{i}_{m}\mathrm{d}x^{m}$ are the unit length local frame one-forms and the $L_{i}$ are the associated radii, expressed in string units ($2\pi\ell_{s}=1)$. The particular example of a seven-dimensional solvmanifold that is of interest here, is realised by the following Mauer-Cartan equations:
$\begin{split}&\mathrm{d}e^{1}=\omega_{a}\;e^{2}\wedge e^{7},\qquad\mathrm{d}e^{2}=-\;\omega_{a}e^{1}\wedge e^{7}\\ &\mathrm{d}e^{3}=\omega_{b}\;e^{4}\wedge e^{7},\qquad\mathrm{d}e^{4}=-\;\omega_{b}e^{3}\wedge e^{7}\\ &\mathrm{d}e^{5}=\omega_{c}\;e^{6}\wedge e^{7},\qquad\mathrm{d}e^{6}=-\;\omega_{c}e^{5}\wedge e^{7}\,.\end{split}$ (3)
These Mauer-Cartan equations can be integrated into the following expressions for the one-forms, in terms of the coordinates $x^{1},...,x^{7}$:
$\begin{split}&e^{1}=\cos\left(\omega_{a}x^{7}\right)\mathrm{d}x^{1}-\sin\left(\omega_{a}x^{7}\right)\mathrm{d}x^{2}\\ &e^{2}=\sin\left(\omega_{a}x^{7}\right)\mathrm{d}x^{1}+\cos\left(\omega_{a}x^{7}\right)\mathrm{d}x^{2}\\ &e^{7}=\mathrm{d}x^{7}\,.\end{split}$ (4)
and for the other directions idem upon the simultaneous cyclic permutation of (1,3,5), (2,4,6) and $(\omega_{a},\omega_{b},\omega_{c})$. This manifold has been considered before, for example in refs. Andriot:2015sia ; Manero2020315 . In essence, this seven-dimensional solvmanifold can be seen as three different copies of the 3d solvmanifold $E_{2}$ (also called ISO(2)) with one common direction (the seventh), see e.g. refs. Grana:2006kf ; Grana:2013ila ; Andriot:2015sia . The three-dimensional space can be divided by different lattices to make it compact, each lattice leading to a different cohomology, see refs. Grana:2013ila ; Andriot:2015sia . However, for this paper, it is enough to take the standard torus lattice $x^{i}\simeq x^{i}+1$, as the Betti numbers will be reduced through an orbifold. Consequentially, this means that the structure constants $\omega_{a,b,c}$ have to be quantised by $2\pi\mathbb{Z}$ when all the $e^{i}$’s have to be preserved by the lattice.
The orbifolds and orientifold
To project out certain forms of the seven-dimensional manifold, and to generate multiple orientifold plane images, it is useful to consider orbifolds of this space. A nice feature of the solvmanifold is that many of the orbifold actions compatible with the flat seven-torus $\mathbb{T}^{7}$, are also compatible here. For example, the standard $\Gamma=\mathbb{Z}_{2}^{3}$ orbifold of the flat torus $\mathbb{T}^{7}$ can be considered here as well. The elements of $\Gamma$ can be written as follows:
$\begin{split}&\alpha(x)=(-x^{1},-x^{2},x^{3},x^{4},-x^{5},-x^{6},x^{7})\\ &\beta(x)=(x^{1},x^{2},-x^{3},-x^{4},-x^{5},-x^{6},x^{7})\\ &\gamma(x)=(-x^{1},x^{2},-x^{3},x^{4},-x^{5},x^{6},-x^{7})\,.\end{split}$ (5)
@@
These three $\mathbb{Z}_{2}$ actions preserve the Mauer-Cartan equations (3), square to one and each have 16 fixed points, just as their combinations $\alpha\beta$, $\gamma\beta$, $\alpha\gamma$, $\alpha\beta\gamma$. For example, the fixed points of $\gamma(x)=x$ are at $x^{1,3,5,7}=\{0,\frac{1}{2}\}$, for all $x^{2,4,6}$. After orbifolding the twisted torus, the space should be orientifolded by an O5-involution for further purposes of this paper, similar to ref. Emelin:2021gzx . Such an involution that also leaves the Mauer-Cartan equations (3) invariant, is the following:
$\sigma(x)=(-x^{1},x^{2},-x^{3},x^{4},-x^{5},x^{6},-x^{7})\,,$ (6)
@@
acting on the coordinates in the same way as $\gamma$. The hypersurfaces of fixed points of this orientifold involution correspond to O5-planes, as they are of codimension 4. Since the involution acts the same as $\gamma$, the orientifold $\sigma$ has the same fixed points as $\gamma$, at $x^{1,3,5,7}=\{0,\frac{1}{2}\}$. Moreover, all orbifold images of $\sigma$, i.e. $\sigma\alpha$, $\sigma{\alpha\beta}$, … have fixed points too, which all lead to different sets of O5-planes. Finally, $\sigma\gamma$ leaves everything invariant and leads to an O9-plane, meaning that one can consider type I string theory as well, after including 32 D9-branes. This orbifold and orientifold were already considered in ref. Emelin:2021gzx , and the inclusion of the O9-plane was useful for matching results with 3d compactifications of heterotic string theory, see e.g. deIaOssa:2019cci ; delaOssa:2021cgd ; delaOssa:2024dzo , due to heterotic - type I string theory duality. Nevertheless, it is not well-understood how the orbifold singularities of such spaces can be resolved.
@@
For this work, it is important to reduce the number of O5-plane sets, and additionally remove the O9-plane so that one remains within type IIB string theory. This can be achieved by changing the orbifold actions by introducing shifts as well. Such types of orbifolds were crucial in Joyce’s construction of compact spaces with $G_{2}$-holonomy Joyce:1996i ; Joyce:1996ii . As a first example, one can consider dividing out by $\hat{\Gamma}=\mathbb{Z}_{2}^{3}$, with the same action $\alpha$, but where shifts are introduced in $\beta$ and $\gamma$:
$\begin{split}&\hat{\alpha}(x)=\biggl{(}-x^{1},-x^{2},x^{3},x^{4},-x^{5},-x^{6},x^{7}\biggr{)}\\ &\hat{\beta}(x)=\biggl{(}x^{1},x^{2},-x^{3},-x^{4},-x^{5},-x^{6},x^{7}+\frac{1}{2}\biggr{)}\\ &\hat{\gamma}(x)=\biggl{(}-x^{1}+\frac{1}{2},x^{2},-x^{3}+\frac{1}{2},x^{4},-x^{5}+\frac{1}{2},x^{6},-x^{7}\biggr{)}\,.\end{split}$ (7)
@@
Only the elements $\hat{\alpha}$ and $\hat{\gamma}$ have fixed points (16 each), all the other combinations do not. Notice that the extra shift in $x^{7}$ requires one to quantise the structure constants $\omega_{a,b,c}$ in $4\pi\mathbb{Z}$ instead of $2\pi\mathbb{Z}$, if one wants all one-forms $e^{i}\to\pm e^{i}$ to just flip sign under all orbifold actions. [^4] Note that a similar quantisation condition also appeared for nilmanifolds, e.g. in ref. Cribiori:2021djm . According to the criteria of Joyce Joyce:1996i ; Joyce:1996ii , this orbifold on the flat $\mathbb{T}^{7}$ has singularities of the ‘good’ type, meaning that they do not intersect and that a blow-up procedure can resolve them. Every fixed point is locally a three-torus on the flat $\mathbb{T}^{7}$. The group element $\hat{\gamma}$ reduces the set of independent fixed $\mathbb{T}^{3}$’s of $\hat{\alpha}$ by half, to 8, hence there are eight singularities of the local form $\mathbb{R}^{4}/\mathbb{Z}_{2}\times\mathbb{T}^{3}$. Similarly, the fixed points of $\hat{\gamma}$ are shuffled around by both $\hat{\alpha}$ and $\hat{\beta}$, leading to only 4 independent singularities. The blow-up procedure generates three local three-cycles for every independent singularity, leading to $(8+4)\times 3=36$ additional three-cycles, whose associated blow-up moduli are said to belong to the twisted sector. There are also 12 local two-cycles from the blow-up. Although it is not straightforward to extrapolate this to the twisted torus $\tilde{\mathbb{T}}^{7}$ here, one could argue that, in a neighbourhood of the fixed points of $\hat{\gamma}$, the twisted torus looks locally like the flat one. For $\hat{\alpha}$ on the other hand, the fixed loci are the three-dimensional solvmanifolds $E_{2}$ instead of $\mathbb{T}^{3}$’s. However, if one would choose to set $\omega_{b}=0$, then they are $\mathbb{T}^{3}$’s, and the arguments above might apply again, although one has to treat this reasoning with caution.
@@
Further orientifolding with the same involution $\sigma$ of eq. (6) results in fixed points of codimension four for $\sigma$, $\sigma\hat{\alpha}$, $\sigma\hat{\beta}$ and $\sigma{\alpha\hat{\beta}}$ , leading to 4 sets of O5-planes. These are summarised in Table 1. The actions $\sigma\hat{\gamma}$, $\sigma{\hat{\alpha}\hat{\gamma}}$, $\sigma{\hat{\beta}\hat{\gamma}}$, $\sigma{\hat{\alpha}\hat{\beta}\hat{\gamma}}$ have no fixed points: for each of them, there is at least one coordinate $x^{i}$ that is mapped into $x^{i}+1/2$, hence no further orientifold images are created. The orientifold involution does not reduce the amount of independent orbifold singularities and therefore does not seem to affect the twisted sector.
$x^{1}$ $x^{2}$ $x^{3}$ $x^{4}$ $x^{5}$ $x^{6}$ $x^{7}$
O5σ -- $\times$ -- $\times$ -- $\times$ --

### v3

# Supersymmetric scale-separated AdS3 orientifold vacua of type IIB

@@ Abstract
I construct supersymmetric AdS3 vacua of type IIB string theory that exhibit parametric scale separation in the controlled regime. These solutions arise from compactifications on seven-dimensional manifolds equipped with co-closed $G_{2}$-structures, in the presence of orientifold planes preserving minimal supersymmetry. I focus on $\mathbb{Z}_{2}^{3}$ orbifolds of a specific solvmanifold and all seven-dimensional nilmanifolds, each requiring distinct configurations of intersecting O5-planes, which are treated in the smeared approximation. Weak string coupling and large volumes can be achieved for classical backgrounds on both nilmanifolds and solvmanifolds by tuning unbounded fluxes to large values. This achieves scale separation in the nilmanifold case, but the same cannot be concluded for the solvmanifolds. Furthermore, for some solutions, the holographic field theory operators dual to the lightest scalar fields have integer conformal dimensions at tree level, as in other scale-separated type IIA orientifold backgrounds.
@@ 1 Introduction
where $L_{\mathrm{KK}}$ and $L_{\mathrm{AdS}}$ are the KK and AdS radius respectively. Some recent progress has provided strong arguments against scale separation in specific settings. For instance, refs. Cribiori:2022trc ; Bobev:2023dwx ; Cribiori:2023ihv ; Perlmutter:2024noo indicate that scale separation for compactifications of several theories with extended supersymmetry is unattainable. Moreover, it has been argued in ref. Gautason:2015tig that for compactifications without sources, scale separation cannot occur for isotropic manifolds of positive scalar curvature, more precisely for when the KK scale decouples from the scalar curvature, and due to the lack of concrete examples it was conjectured not to happen in ref. Collins:2022nux . All these findings support the strong version of the Anti-de Sitter distance conjecture Lust:2019zwm , implying that no supersymmetric scale-separated vacua exist. However, the minimally supersymmetric theories with sources are not constrained yet by the arguments of refs. Cribiori:2022trc ; Bobev:2023dwx ; Cribiori:2023ihv ; Perlmutter:2024noo ; Gautason:2015tig ; Collins:2022nux , and indeed, there are a few scale-separated models in string theory, therefore violating the Anti-de Sitter distance conjecture Lust:2019zwm . Nevertheless, their validity is questioned and is an active research topic in the context of the swampland program. For a recent review on this issue, see ref. Coudarchet:2023mfs .
@@
The current state of the art of supersymmetric scale-separated vacua can be summarised as a type IIA vs. IIB dichotomy: in type IIA, there are scale-separated vacua where all moduli are stabilised using fluxes and orientifold planes, hence they are referred to as classical vacua. The separation of scales is achieved by tuning unbounded, integer-quantised fluxes to infinity. Examples in four dimensions are compactifications on manifolds with SU(3)-structure, usually toroidal orbifolds, e.g. the DGKT solutions in massive IIA DeWolfe:2005uu ; Camara:2005dc and analogue solutions in massless IIA Cribiori:2021djm ; Carrasco:2023hta , of which some can be lifted to solutions of M-theory which appear to have weak $G_{2}$-structure VanHemelryck:2024bas . Examples in three dimensions are compactifications on manifolds with $G_{2}$-holonomy, as considered in Farakos:2020phe ; VanHemelryck:2022ynr . [^1] See also refs. Farakos:2023nms ; Farakos:2023wps ; Tringas:2023vzn for variations on these setups, and ref. Andriot:2025cyi for a discussion on scale separation in rolling solutions. Although the holographic dual field theories for these models are not known, refs. Conlon:2021cjk ; Apers:2022tfm ; Quirant:2022fpn found, quite surprisingly, that the mass spectrum of the scalars in the four-dimensional scale-separated vacua is such that the putative holographic dual operators have integer conformal dimensions. [^2] This feature is expected to break down by small one-loop corrections, see ref. Plauschinn:2022ztd . This feature does not appear in the three-dimensional vacua of ref. Farakos:2020phe , as shown in ref. Apers:2022zjx .
@@
The situation in type IIB is quite different. Non-perturbative effects are usually necessary for full moduli stabilisation and obtaining scale separation. Examples are four-dimensional Calabi-Yau compactifications with three-form fluxes and a gaugino condensate or Euclidean D3-instantons, such as in the KKLT and LVS scenarios Kachru:2003aw ; Conlon:2005ki ; Balasubramanian:2005zx . Recent progress has been made in finding solutions with a large hierarchy of scales (i.e. finding a small flux superpotential), see for example refs. Demirtas:2019sip ; Demirtas:2020ffz ; Demirtas:2021nlu ; Demirtas:2021ote ; McAllister:2024lnt , however, these scenarios have been criticised on different fronts as well, see e.g. refs. Bena:2018fqc ; Gao:2020xqh ; Bena:2020xrh ; Gao:2022fdi ; Gao:2022uop ; Lust:2022lfc for the most recent criticism. So far, no supersymmetric and scale-separated vacua of type IIB have been found where only classical ingredients stabilise all moduli, while maintaining weak coupling and large volume. [^3] The 4d vacua of refs. Caviezel:2009tu ; Petrini:2013ika are not well-controlled, as some one-cycles are becoming small in string units, see ref. Cribiori:2021djm . Recently, the possibility of scale separation in type IIB orientifold compactifications on flat manifolds has been studied in Tringas:2025uyg .
@@
However, this paper presents the first such examples. These are realised by compactifying type IIB string theory on seven-dimensional manifolds equipped with a co-closed $G_{2}$-structure. Compactifications on such backgrounds have been considered before in refs. Passias:2020ubv ; Emelin:2021gzx and VanHemelryck:2022ynr . In particular, this paper studies $\mathbb{Z}_{2}^{3}$ orbifolds of two main classes of group manifolds: solvmanifolds and nilmanifolds. For the former, a specific solvmanifold is considered, which has been described in the literature before, see e.g. refs. DallAgata:2005zlf ; Andriot:2015sia ; Manero2020315 . After orientifolding these specific spaces with an O5-involution, one obtains one class of manifolds with four (intersecting) sets of (smeared) O5-planes and another class with only one set of O5-planes. The former is reminiscent of the type IIA vacua with 4 sets of (intersecting) O6-planes, whereas the latter case can be regarded as the supersymmetric cousin of a scale-separated, but non-supersymmetric type IIB vacuum found in ref. Arboleya:2024vnp ; Arboleya:2025ocb , which are related through skew-whiffing. These solutions achieve large cycle volumes, weak coupling and scale-separation between the AdS scale and cycle volumes, but those do not seem to provide a good estimate of the KK scale. The second class of manifolds comprises nilmanifolds, for which the latter issue does not arise. The study of seven-dimensional nilmanifolds that admit a co-closed $G_{2}$-structure has received great attention recently, e.g. in refs. MR2811660 ; MR3739330 ; MR4626831 ; MR4789076 . There are only 12 nilmanifolds consistent with the orbifold $\mathbb{Z}_{3}^{2}$, and scale-separated solutions in the controlled regime were found in this paper on only 6 of them. The simplest nilmanifold, being a direct product of the 3d Heisenberg nilmanifold and a flat four-torus, exhibits integer conformal dimensions for the operators dual to the scalars, all the others do not.
The analysis here uses the tools of refs. Emelin:2021gzx ; Passias:2020ubv ; VanHemelryck:2022ynr , but is different in the sense that in those references, models with seven (intersecting) sets of O5-planes were considered instead of fewer. Particularly in ref. Emelin:2021gzx , it was argued that scale separation can be achieved in their setup if one includes additional D5-branes to cancel some of the tadpoles, and if one can make the curvature small compared to the compactification radius. However, the latter clashes with geometric flux quantisation when the manifold is isotropic. The setups of this paper circumvent these issues in the sense that the manifolds are non-isotropic, and one also does not necessarily need wrapped D5-branes due to the choice of orbifold generators.
This paper is organised as follows: section 2 discusses the $\mathbb{Z}_{2}^{3}$ orbifold and orientifold of seven-dimensional twisted tori and their relation to co-closed $G_{2}$-structures. Section 3 briefly reviews type IIB compactifications on manifolds with $G_{2}$-structure, and discusses the specific compactifications on some solvmanifolds and nilmanifolds. Finally, the results are summarised in section 4.
2 $\mathbb{Z}_{2}^{3}$ orbifolds of twisted tori and co-closed $G_{2}$-structures
@@
Before discussing three-dimensional flux compactifications of type IIB string theory, it is useful to focus on the geometries and especially the orbifolds that are used for this purpose. Especially nil- and solvmanifolds play a special role here, which are group manifolds arising from a nilpotent or solvable group, respectively. They can be regarded as a twisted seven-torus, denoted by $\tilde{\mathbb{T}}^{7}$, for which the metric in the vielbein parametrisation is
@@
where the $e^{i}=e^{i}_{m}\mathrm{d}x^{m}$ are the unit length local coframe one-forms and the $L_{i}$ are the associated radii, expressed in string units ($2\pi\ell_{s}=1)$. These one-forms are not necessarily closed, and indeed, they satisfy the Maurer-Cartan equations:
$\mathrm{d}e^{a}=\frac{1}{2}f^{a}{}_{{}{bc}}\;e^{b}\wedge e^{c}\,$, (3)
where the $f^{a}{}_{{}{bc}}$ are the structure constants of the Lie algebra underlying the group manifold. To project out certain forms of the seven-dimensional manifold and to generate multiple orientifold plane images, it is useful to consider orbifolds of this space. In this work, the $\Gamma=\mathbb{Z}_{2}^{3}$ orbifold is considered, very similar to refs. DallAgata:2005zlf ; Emelin:2021gzx . The three $\mathbb{Z}_{2}$ generators $\alpha$, $\beta$ and $\gamma$ act on the coframe one-forms as follows:
$\begin{split}&\alpha(e)=(-e^{1},-e^{2},e^{3},e^{4},-e^{5},-e^{6},e^{7})\,,\\ &\beta(e)=(e^{1},e^{2},-e^{3},-e^{4},-e^{5},-e^{6},e^{7})\,,\\ &\gamma(e)=(-e^{1},e^{2},-e^{3},e^{4},-e^{5},e^{6},-e^{7})\,.\end{split}$ (4)
The Maurer-Cartan equations must be invariant under the orbifold actions, which do not allow for all structure constants, leaving only the following ones:
$f^{5}{}_{14},\quad f^{5}{}_{23},\quad f^{6}{}_{13},\quad f^{6}{}_{24},\quad f^{7}{}_{12},\quad f^{7}{}_{34},\quad f^{7}{}_{56}\quad+\;\text{permutations}\,.$ (5)
To investigate the number of fixed points of the orbifold and orientifold later, one needs to work in a coordinate representation of the metric and coframe one-forms, with coordinates $x^{1},...,x^{7}$. It is instructive to consider the standard $\Gamma=\mathbb{Z}_{2}^{3}$ orbifold of the ordinary torus $\mathbb{T}^{7}$ first, as it will act similarly on the coordinate systems of the solv- and nilmanifolds. The elements of $\Gamma$ can be written as follows:
$\begin{split}&\alpha(x)=(-x^{1},-x^{2},x^{3},x^{4},-x^{5},-x^{6},x^{7})\,,\\ &\beta(x)=(x^{1},x^{2},-x^{3},-x^{4},-x^{5},-x^{6},x^{7})\,,\\ &\gamma(x)=(-x^{1},x^{2},-x^{3},x^{4},-x^{5},x^{6},-x^{7})\,.\end{split}$ (6)
@@
These three $\mathbb{Z}_{2}$ actions square to one and each have 16 fixed points, just as their combinations $\alpha\beta$, $\gamma\beta$, $\alpha\gamma$, $\alpha\beta\gamma$. For example, the fixed points of $\gamma(x)=x$ are at $x^{1,3,5,7}=\{0,\frac{1}{2}\}$, for all $x^{2,4,6}$. After orbifolding, one should orientifold the space by an O5-involution for further purposes of this paper, similar to ref. Emelin:2021gzx . Such an involution that also leaves the Maurer-Cartan equations (3) invariant, is for instance:
$\sigma_{\gamma}(x)=(-x^{1},x^{2},-x^{3},x^{4},-x^{5},x^{6},-x^{7})\,$, (7)
@@
acting on the coordinates in the same way as $\gamma$. The hypersurfaces of fixed points of this orientifold involution correspond to O5-planes, as they are of codimension 4. Since the involution acts the same as $\gamma$, the orientifold $\sigma_{\gamma}$ has the same fixed points as $\gamma$, at $x^{1,3,5,7}=\{0,\frac{1}{2}\}$. Moreover, all orbifold images of $\sigma_{\gamma}$, i.e. $\sigma_{\gamma}\alpha$, $\sigma_{\gamma}{\alpha\beta}$, … have fixed points too, which all lead to seven different (intersecting) sets of O5-planes. Finally, $\sigma_{\gamma}\gamma$ leaves everything invariant and leads to an O9-plane, meaning that one can consider type I string theory as well, after including 32 D9-branes. This orbifold and orientifold were already considered in ref. Emelin:2021gzx , and the inclusion of the O9-plane was useful for matching results with 3d compactifications of heterotic string theory, see e.g. deIaOssa:2019cci ; delaOssa:2021cgd ; delaOssa:2024dzo , due to heterotic - type I string theory duality. Nevertheless, it is not well-understood how the orbifold singularities of such spaces can be resolved.
@@
For this work, it is important to reduce the number of intersecting O5-plane sets and additionally remove the O9-plane so that one remains within type IIB string theory. This is achieved by including shifts in the orbifold actions as well. Such types of orbifolds were crucial in Joyce’s construction of compact spaces with $G_{2}$-holonomy Joyce:1996i ; Joyce:1996ii . This can be achieved by adding 7-vectors $\vec{a}$, $\vec{b}$ and $\vec{c}$ whose entries are either 0 or $1/2$:
$\begin{split}&\hat{\alpha}(x)=\biggl(-x^{1},-x^{2},x^{3},x^{4},-x^{5},-x^{6},x^{7}\biggr)+\vec{a}\,,\\ &\hat{\beta}(x)=\biggl(x^{1},x^{2},-x^{3},-x^{4},-x^{5},-x^{6},x^{7}\biggr)+\vec{b}\,,\\ &\hat{\gamma}(x)=\biggl(-x^{1},x^{2},-x^{3},x^{4},-x^{5},x^{6},-x^{7}\biggr)+\vec{c}\,.\end{split}$ (8)
As a first example, which is relevant for the solvmanifold, one should consider the orbifold $b_{7}=c_{1}=c_{3}=c_{5}=1/2$ and all the rest are vanishing.
$\hat{\Gamma}_{4}\;:\qquad b_{7}=c_{1}=c_{3}=c_{5}=1/2\,.$ (9)
@@
For this orbifold, only the elements $\hat{\alpha}$ and $\hat{\gamma}$ have fixed points (16 each), all the other combinations do not. According to the criteria of Joyce Joyce:1996i ; Joyce:1996ii , this orbifold on the ordinary $\mathbb{T}^{7}$ has singularities of the ‘good’ type, meaning that they do not intersect and that a blow-up procedure can resolve them. Every fixed point is locally a three-torus on the ordinary $\mathbb{T}^{7}$. The group element $\hat{\gamma}$ reduces the set of independent fixed $\mathbb{T}^{3}$’s of $\hat{\alpha}$ by half, to 8, hence there are eight singularities of the local form $\mathbb{R}^{4}/\mathbb{Z}_{2}\times\mathbb{T}^{3}$. Similarly, the fixed points of $\hat{\gamma}$ are shuffled around by both $\hat{\alpha}$ and $\hat{\beta}$, leading to only 4 independent singularities. The blow-up procedure generates three local three-cycles for every independent singularity, leading to $(8+4)\times 3=36$ additional three-cycles, whose associated blow-up moduli are said to belong to the twisted sector. There are also 12 local two-cycles from the blow-up.
@@
Further orientifolding with the same involution $\sigma_{\gamma}$ of eq. (7) results in fixed points of codimension four for $\sigma_{\gamma}$, $\sigma_{\gamma}\hat{\alpha}$, $\sigma\hat{\beta}$ and $\sigma_{\gamma}{\hat{\alpha}\hat{\beta}}$ , leading to 4 sets of O5-planes. These are summarised in Table 1. The actions $\sigma_{\gamma}\hat{\gamma}$, $\sigma_{\gamma}{\hat{\alpha}\hat{\gamma}}$, $\sigma_{\gamma}{\hat{\beta}\hat{\gamma}}$, $\sigma_{\gamma}{\hat{\alpha}\hat{\beta}\hat{\gamma}}$ have no fixed points: for each of them, there is at least one coordinate $x^{i}$ that is mapped into $x^{i}+1/2$, hence no further orientifold images are created, neither does the orientifold involution reduce the amount of independent orbifold fixed points. The orientifold planes are summarised in Table 1.
$\hat{\Gamma}_{4}$ $x^{1}$ $x^{2}$ $x^{3}$ $x^{4}$ $x^{5}$ $x^{6}$ $x^{7}$